OpenStudy (anonymous):

Suppose that a car normally sells for $16,500 and is now on sale for $14,000. What is the percent of discount? If necessary, round your answer to the nearest tenth.

OpenStudy (anonymous):

let the normal price be x let the new price be y let the discount be d discount(d) = normal price(x) - new price(y)' d = x - y d = 16500 - 14000 d = 2500 Now we have d solving the percent discount should be easy. Let me know if you cant solve it.

OpenStudy (anonymous):

wellits bettween %3 or %20 or %25 or %80